# HREB-CRF: Hierarchical Reduced-bias EMA for Chinese Named Entity Recognition

### v1

@@ Abstract
Incorrect boundary division, complex semantic representation, and differences in pronunciation and meaning often lead to errors in Chinese Named Entity Recognition(CNER). To address these issues, this paper proposes HREB-CRF framework: Hierarchical Reduced-bias EMA with CRF. The proposed method amplifies word boundaries and pools long text gradients through exponentially fixed-bias weighted average of local and global hierarchical attention. Experimental results on the MSRA, Resume, and Weibo datasets show excellent in F1, outperforming the baseline model by 1.1%, 1.6%, and 9.8%. The significant improvement in F1 shows evidences of strong effectiveness and robustness of approach in CNER tasks.
Equal contribution  Corresponding author
Keywords NER  $\cdot$ Hierarchical Attention  $\cdot$ RoBERTa  $\cdot$ Token Classification
1 Introduction
@@
Therefore, inspired by the MEGA structure[5], a Reduced-biased Hierarchical EMAttention framework (RHEMA) is proposed integrated with RoBERTa[6]. This hierarchical design, shown as Figure-1 ensures both short-range (local) and long-range (global) dependencies are jointly modeled, addressing segmentation ambiguities. The reduced-bias connection further mitigates the risk of gradient vanishing or exploding when stacking deep layers, thus providing more stable parameter updates.
Figure 1: Constructing on the uniqueness of the decay weight mechanism in the exponential averaging method, a reduced-biased exponential averaging layer is incorporated into the attention mechanism.
To sum up, proposed contributions are mainly as follows:
@@
  3. 3.
2 Related Work
Deep learning methods have significant advantages in named entity recognition tasks. One of its main advantages is the ability to capture contextual information by learning nested representations, helping to identify the boundaries of entities more accurately. Deep learning models can automatically learn abstract features in text without manually designing features, thereby more comprehensively understanding the multi-faceted information of entities, including morphology, syntax, and semantics. In addition, deep learning models can handle long-distance dependencies, capture dependencies at different distances in text through operations such as recursion or convolution, and improve the understanding of the entire text structure, thus improving the accuracy of named entity recognition. And it can also be combined with conditional random fields (CRF) to globally model the label sequence, consider the joint probability of the overall sequence, ensure the consistency between labels, and avoid unreasonable label sequences. This integration method makes the model better in NER. More comprehensive and accurate identification of entities in tasks.
@@ 2 Related Work
Qiu et al. [7] highlighted that discrepancies in text semantics caused by inaccurate tokenization are a distinctive challenge in Chinese named entity recognition (NER) compared to English NER. Different tokenization schemes can lead to varying evaluation metrics. He et al.’s study [8] found that character-based segmentation in Chinese NER outperforms word-based segmentation, indicating that a finer granularity is advantageous for Chinese language processing. Further notable findings show that static embedding techniques, such as Word2Vec [9], while effective in deep neural networks, fail to address polysemy—a characteristic unique to Chinese language processing. Building on this, the CWVF-BiLSTM-CRF model developed by Ye et al. [10] demonstrates that incorporating both character and word inputs can reduce accuracy errors caused by tokenization while improving entity recognition precision.
Cao et al. [11] uses an adversarial transfer learning model to predict text boundary intervals, which effectively filters noise within sentences. Lattice-LSTM [1] is a pioneering method in Chinese NER, as prior research consistently shows that dictionary-based approaches can significantly enhance Chinese NER performance, with many studies adapting their models based on this algorithm. To address lexical conflicts, Gui et al. [12] introduced the LR-CNN model, which employs CNN layers for feature extraction and integrates attention mechanisms between CNN layers to manage vocabulary conflicts through a "Rethinking" mechanism. Additionally, Liu’s work [13] encodes word information using multiple representations, converting words into fixed-length feature vectors. Lastly, the FLAT-Transformer model by Li et al. [2] leverages a lattice-based structure to represent text, converting character grids into a flattened representation that enhances NER performance.
Moreover, specialized data structures can effectively capture named entity features. The Tree-LSTM model [14] demonstrates that using a topological network structure within data structures can not only integrate sub-feature information but also provide richer topological representations. DyLex [15], proposed by Wang et al., is a BERT-based dictionary augmentation method that addresses noise in text sequences under supervised training and allows dictionary updates without requiring retraining. Liu [16] introduced the TFM network, aimed at enhancing character understanding by integrating positional information, thereby capturing both relative and absolute positional relationships among characters within the model to improve character interrelation comprehension.
Besides, many studies also explore Chinese word structure, phonetics, and radicals as features for named entity recognition. Mai [17] examines different practical features of homophones based on lexical and phonetic characteristics. Li [18] leverages syntactic dependency trees to guide global attention mechanisms through rule-based methods.
3 Methodology
3.1 NER Task Definition
The NER task is a typical sequence labeling task. Based on a $token$ sequence $x={x_{1},x_{2},\cdots,x_{n}}$ of a given length $n$, the output $y={y_{1},y_{2},\cdots,y_{n}}$ is predicted. Where $x_{i}\in\text{T}$, $T$ represents the vocabulary. In this study, the BIO labeling method is used to label the category $y$. The entire task flow requires the construction of a text processing sequence, as shown in the Equation-1.
$input$ =$=\{x_{1},x_{2},\cdots,x_{n}\}$ Tokenization (1)
$vectoried$ =$=\{\text{i}_{1},\text{i}_{2},\cdots,\text{i}_{n}\}$ BERT Embedding
$featured$ =$=\{\text{v}_{1},\text{v}_{2},\cdots,\text{v}_{n}\}$ Feature Extraction
$cls$ =$=\{\text{f}^{1}_{1},\text{f}^{1}_{2},\cdots,\text{f}^{C}_{n}\}$ Segmentation Weight
$y$ =$=\{c_{1},c_{2},\cdots,c_{n}\}$ Sequence Dependency
where $C$ denotes number of classes.
In detail, the whole process starts with the input text sequence, and the text is converted into a set of $token$ through word segmentation. Then, these $token$ are mapped to the vector space through the BERT embedding layer to generate an embedded representation of each $token$. Subsequently, the embedded representation is passed to the feature extraction module to obtain deeper contextual features. The classification layer calculates multi-category scores for the feature vectors at each position to obtain the weights corresponding to each category. Finally, CRF layer is used to further model the dependencies between sequences, and the argmax operation is used to select the label sequence with the highest probability to obtain the final prediction output.
3.2 Overview of HREB-CRF Architecture
@@
Proposed model is shown in Figure-2. BERT structure model is used as the pre-training head, and output the start and end (B-Cls and last I-Cls structure annotation) features of the CRF-constrained entity. On this basis, the additional Hierarchical EMA module is used to enhance capturing the short-distance relationship of the text vector. Compared with traditional attention method, this improved method has the ability to extract both short-range and long-range features. Next, this hierarchical dependency weights are input into the Bi-LSTM layer to capture the context and strengthen the long-distance dependency capabilities, and finally input into the CRF for sequence dependency decoding. In addition, a reduced-biased module is added to each layer to reduce the possibility of gradient disappearance.
@@
3.3 Exponentially Moving Average
Moving average method is widely used in sequence analysis tasks, such as stock and anomaly detection. It is proposed to reduce the global impact of local outliers anomalies. Generally speaking, most of the usage of MA are: simple moving average(Equation-2), weighted moving average(Equation-3), cumulative moving average(Equation-4) and exponentially weighted moving average(Equation-5).
@@ 3.3 Exponentially Moving Average
In the exponentially moving average (EM) formula, $\alpha$ is used as a weighting factor to determine the degree of attenuation of historical data. A larger $\alpha$ value indicates that current data is given more importance and the reliance on earlier data is weakened. Therefore, EM can effectively reduce the impact of random factors in the sequence on the overall trend, because as time goes by, $(1-\alpha)^{t}\cdot y_{t}$ will gradually decrease and tend to zero, gradually diluting the weight of earlier data.
@@
3.4 HEMA
Since the exponential averaging method is unique in the decay weight mechanism, we add the reduced-biased exponential averaging layer to the attention mechanism. Based on the MEGA method of [5], reduced-biased Hierarchical EMA is introduced as improved model. This improved architecture is shown in the Figure-3.
Figure 3: Proposed RHEMA architecture with added reduced-biased module. The figure illustrates the architecture of the RHEMA framework for text processing, enhanced by the addition of a reduced-bias module, marked in red. The input text is initially transformed through a token embedding layer. The core RHEMA block includes sequential processing through Batch Normalization and a FeedForward network. A novel "Reduced-Bias Add" module, shown with red arrows, is integrated to mitigate potential bias in the output, while preserving a feedback loop to earlier stages. This enhancement aims to improve robustness and fairness in predictions. The final output weights are generated post these computations.
RHEMA uses GAU[19] to share feature dimensions. First, pass it to the EMA layer at Equation-6, then use the SiLU self-gated activation function to linearly project to z dimension, and add nonlinear features.
@@ 3.4 HEMA
With the advantage of SiLU, the operation here can compress the originally redundant d-dim input into a more compact z-dim representation while retaining the context information. Equation-8 represents the specific calculation process of SiLU.
@@
Then, through the output of Equation-7, Query(Equation-9), Key(Equation-10) and Value(Equation-11) is calculated. Among them, the bias terms ($\mu_{q}$, $\mu_{k}$ and $b_{v}$) add additional degrees of freedom to the attention mechanism, which allows the query and key to adjust their respective baseline values based on the shared representation $Z$ to make them more suitable for broader tasks. And $v$ refers to the extension dim of value series.
@@
Next, the result of the key-value pair query can be expressed as Equation-12.
@@
by simulating the gating mechanism of GRU, activation output(Equation-13) is introduced. It affects the influence of each cell on the global affection.
@@
Finally, the output is represented as Equation-14.
@@
The single-head attention calculation method mentioned in Equation-14 can be expressed as Figure-4.
@@
The output $Y$ combines the candidate activation state $\hat{Y}$ and the original input $X_{in}$, and the update gate $\Phi$ controls the information fusion ratio between the two parameters. When calculating the candidate activation state $\hat{H}$, the exponential average attention mechanism uses the Laplace attention function(Equation-15) as the core component. Based on the concept of reduced-bias, it can be improved the reduced-bias Laplace function as Equation-16 to adapt to the introduction of deviation inside.
@@
3.5 Reduced-bias Module
@@
Classic resnet can be expressed as Equation-17.
@@
Among Equation-17, dynamic adaptation mechanism parameters $\alpha$ and $\beta$ is introduced. The new network can be expressed as Equation-18.
@@
By deriving Equation-19, $\mathbf{y}_{l}$ can be got by Equation-20.
@@
3.6 Robustly Optimized BERT
@@
4 Experiments
4.1 Datasets Description
@@
The specific properties of the original datasets are shown in the Table-1. It should be noted that since MSRA does not have a validation set by default, all experiments directly use the test-set as the valid-set. For the other two datasets, a split rate of 7:1.5:1.5 is used to evaluate.
Table 1: Properties of the three datasets used.
@@
train 1350 46364 3821
test 270 4365 477
valid 269 4365 463
@@ 4.3 Evaluation Metrics
The negative log-likelihood function, shown as Equation-21 is considered as the loss function to quantitatively evaluate the performance of the model.
@@
4.4 Experimental Results
4.4.1 Comparison Studies
The experimental results on MSRA, Resume and Weibo for comparison studies are shown in the several tables respectively.
Table 2: Comparison Experimental results on MSRA.
@@ 4.4.1 Comparison Studies
Compared with multiple existing methods on MSRA dataset by Table-2, he classic Lattice LSTM model achieves an F1 score of 93.18, demonstrating its effectiveness by combining character and word-level information. However, proposed method surpasses this benchmark with an F1 score of 94.28, demonstrating stronger semantic modeling capabilities. Compared with the 82.62 F1 score of BiLSTM+CRF as the baseline model, proposed structure shows significant advantages in deep feature extraction and context information utilization.
@@
Table 3: Comparison Experimental results at Resume.
@@
Although proposed model has achieved meaningful recall (96.61) and F1 score (96.07) on the Resume dataset, detailed in Table-3. There are still some shortcomings that deserve attention. First, compared with the DAE model, the precision is slightly lower (95.53 vs. 96.92), indicating that there is space for improvement in the prediction accuracy of specific entities.
Table 4: Comparison Experimental results at Weibo.
@@
FLAT Li 2020[2] 63.42
FLAT+BERT Li 2020[2] 68.55
@@
Proposed method achieves an F1 score of 68.60, slightly higher than FLAT+BERT (68.55) and close to Glyce (67.60), demonstrating robustness in complex text processing on Weibo, shown as Table-4. Furthermore, proposed method performs well in recall (70.76), indicating its advantage in identifying more potential entities. However, compared with the precision of DAE (69.68), proposed method still has a certain gap in accuracy, which suggests that proposed model still has room for improvement in accurate prediction when dealing with noisy social media data.
4.4.2 Ablation Studies
In the ablation experiment, shown as Table-5, two experiments are designed to evaluate the real impacts of proposed structure. First, each module is removed to confirm the synergy between different modules. Second, BERT is replaced or removed to compare the impact of word embedding weights on the next layer input.
@@ 4.4.2 Ablation Studies
The experimental results of comparing the base model (S.N.1) with the models that introduce different BERT variants (S.N.2 and S.N.3) show that RoBERTa outperforms BERT in terms of vocabulary vectorization. This shows that RoBERTa has a stronger ability to represent text semantics and can capture contextual information more effectively.
@@
Table 5: Ablation Studies on MSRA
@@ A H R BERT P R F1
8
$\surd$ $\surd$ RoBERTa 94.58 93.98 94.28
4.5 Parameter Selection
@@
Table 6: Effects of different parameters on performance in RHEMA.
@@
As seen from the Table-6, too few attention heads with 16 or 32, may cause the model to be unable to fully capture feature representations, thereby limiting the performance of the model. When the number of attention heads increases to 64, the performance of the model reaches its peak. However, when the number of attention heads is further increased to 128), the performance of the model does not continue to improve, but instead shows a slight decline. This phenomenon may be due to the fact that too many attention heads increase the complexity of the model, thereby introducing the risk of overfitting.
4.6 Case Analysis

### v2

@@ Abstract
Incorrect boundary division, complex semantic representation, and differences in pronunciation and meaning often lead to errors in Chinese Named Entity Recognition (CNER). To address these issues, this paper proposes HREB-CRF framework: Hierarchical Reduced-bias EMA with CRF. The proposed method amplifies word boundaries and pools long text gradients through exponentially fixed-bias weighted average of local and global hierarchical attention. Experimental results on the MSRA, Resume, and Weibo datasets show excellent in F1, outperforming the baseline model by 1.1%, 1.6%, and 9.8%. The significant improvement in F1 shows evidences of strong effectiveness and robustness of approach in CNER tasks. Code available at https://github.com/StanleySun233/HREB-CRF.
Index Terms:
NER, Hierarchical Attention, RoBERTa, Token Classification
I Introduction
@@
Therefore, inspired by the MEGA structure[5], a Reduced-biased Hierarchical EMAttention framework (RHEMA) is proposed integrated with RoBERTa[6]. This hierarchical design, shown as Figure 1 ensures both short-range (local) and long-range (global) dependencies are jointly modeled, addressing segmentation ambiguities. The reduced-bias connection further mitigates the risk of gradient vanishing or exploding when stacking deep layers, thus providing more stable parameter updates.
The contributions of this work can be mainly summarized as the follows:
@@
II Related Work
Deep learning performs well in NER by capturing contextual information, learning abstract features automatically, and handling long-distance dependencies through recursive or convolutional operations. Integrating conditional random fields (CRF) further enhances performance by modeling label sequences globally, ensuring consistency and improving accuracy.
@@ II Related Work
Chinese NER faces unique challenges due to tokenization-induced semantic discrepancies [7]. Character-based segmentation outperforms word-based approaches [8], and while Word2Vec [9] struggles with polysemy, hybrid models like CWVF-BiLSTM-CRF improve accuracy by combining character and word inputs [10]. Adversarial transfer learning aids boundary detection [11], while Lattice-LSTM [1] and its variants (e.g., LR-CNN [12]) leverage dictionaries and attention to resolve lexical conflicts. Multi-representation encoding [13] and FLAT-Transformer’s lattice structures [2] further advance performance.
Specialized data structures enhance NER feature capture. Tree-LSTM [14] integrates sub-features via topological networks, while DyLex [15] augments BERT with updatable dictionaries. Liu et al.’s TFM [16] improves character understanding through positional encoding. Besides, other works on exploiting Chinese linguistic features. Mai et al.[17] studies homophones, and Li et al. [18] uses syntactic dependency trees for global attention.
III Methodology
III-A Overview of HREB-CRF Architecture
@@
Proposed model is shown in Figure 1. BERT structure model is used as the pre-training head, and output the start and end (B-Cls and last I-Cls structure annotation) features of the CRF-constrained entity. On this basis, the additional Hierarchical EMA module is used to enhance capturing the short-distance relationship of the text vector. Compared with traditional attention method, this improved method has the ability to extract both short-range and long-range features. Next, this hierarchical dependency weights are input into the Bi-LSTM layer to capture the context and strengthen the long-distance dependency capabilities, and finally input into the CRF for sequence dependency decoding. In addition, a reduced-biased module is added to each layer to reduce the possibility of gradient disappearance.
@@
III-B Exponentially Moving Average
Moving average method is widely used in sequence analysis tasks, such as stock and anomaly detection. It is proposed to reduce the global impact of local outliers anomalies. Generally speaking, most of the usage of MA are: simple moving average in Equation 1, weighted moving average in Equation 2, cumulative moving average in Equation 3 and exponentially weighted moving average in Equation 4.
@@ III-B Exponentially Moving Average
In the exponentially moving average (EM) formula, $\alpha$ is used as a weighting factor to determine the degree of attenuation of historical data. A larger $\alpha$ value indicates that current data is given more importance and the reliance on earlier data is weakened. Therefore, EM can effectively reduce the impact of random factors in the sequence on the overall trend, because as shift window goes by, $(1-\alpha)^{t}\cdot y_{t}$ will gradually decrease and tend to zero, gradually diluting the weight of earlier data.
@@
III-C HEMA
Since the exponential averaging method is unique in the decay weight mechanism, we add the reduced-biased exponential averaging layer to the attention mechanism. Based on the MEGA method of [5], reduced-biased Hierarchical EMA is introduced as improved model. This improved architecture is shown in the Figure 2.
Figure 2: Proposed RHEMA architecture with added reduced-biased module. The figure illustrates the architecture of the RHEMA framework for text processing, enhanced by the addition of a reduced-bias module, marked in red. The core RHEMA block includes sequential processing through Batch Normalization and a FeedForward network. A novel ”Reduced-Bias Add” module, shown with red arrows, is integrated to mitigate potential bias in the output, while preserving a feedback loop to earlier stages.
RHEMA uses GAU[19] to share feature dimensions. First, pass it to the EMA layer at Equation 5, then use the SiLU self-gated activation function to linearly project to z dimension, and add nonlinear features.
@@ III-C HEMA
With the advantage of SiLU, the operation here can compress the originally redundant d-dim input into a more compact z-dim representation while retaining the context information. Equation 7 represents the specific calculation process of SiLU.
@@
Then, through the output of Equation 6, Query in Equation 8, Key in Equation 9 and Value in Equation 10 is calculated. Among them, the bias terms ($\mu_{q}$, $\mu_{k}$ and $b_{v}$) add additional degrees of freedom to the attention mechanism, which allows the query and key to adjust their respective baseline values based on the shared representation $Z$ to make them more suitable for broader tasks. And $v$ refers to the extension dim of value series.
@@
Next, the result of the key-value pair query can be expressed as Equation 11.
@@
by simulating the gating mechanism of GRU, activation output in Equation 12 is introduced. It affects the influence of each cell on the global affection.
@@
Finally, the output is represented as Equation 13.
@@
The single-head attention calculation method mentioned in Equation 13 can be expressed as Figure 3.
@@
The output $Y$ combines the candidate activation state $\hat{Y}$ and the original input $X_{in}$, and the update gate $\Phi$ controls the information fusion ratio between the two parameters. When calculating the candidate activation state $\hat{H}$, the exponential average attention mechanism uses the Laplace attention function in Equation 14 as the core component. Based on the concept of reduced-bias, it can be improved the reduced-bias Laplace function as Equation 15 to adapt to the introduction of deviation inside.
@@
III-D Reduced-bias Module
@@
Classic resnet can be expressed as Equation 16.
@@
Among Equation 16, dynamic adaptation mechanism parameters $\alpha$ and $\beta$ is introduced. The new network can be expressed as Equation 17.
@@
By deriving Equation 18, $\mathbf{y}_{l}$ can be got by Equation 19.
@@
III-E Robustly Optimized BERT
@@
IV Experiments
IV-A Datasets Description
@@
The specific properties of the original datasets are shown in the Table I. It should be noted that since MSRA does not have a validation set by default, all experiments directly use the test-set as the valid-set. For the other two datasets, a split rate of 70%:15%:15% is used to evaluate.
TABLE I: Properties of the three datasets used.
@@
Train 1,350 46,364 3,821
Test 270 4,365 477
Valid 269 4,365 463
@@ IV-C Evaluation Metrics
The negative log-likelihood function, shown as Equation 20 is considered as the loss function to quantitatively evaluate the performance of the model.
@@
IV-D Experimental Results
IV-D1 Comparison Studies
The experimental results on MSRA, Resume and Weibo for comparison studies are shown in the several tables.
TABLE II: Comparison Experimental results on MSRA.
@@ IV-D1 Comparison Studies
Compared with multiple existing methods on MSRA dataset by Table II, the classic Lattice LSTM model achieves an F1 score of 93.18, demonstrating its effectiveness by combining character and word-level information. However, proposed method surpasses this benchmark with an F1 score of 94.28, demonstrating stronger semantic modeling capabilities. Compared with the 82.62 F1 score of BiLSTM+CRF as the baseline model, proposed structure shows significant advantages in deep feature extraction and context information utilization.
@@
TABLE III: Comparison Experimental results at Resume.
@@
Although proposed model has achieved meaningful recall (96.61) and F1 score (96.07) on the Resume dataset, detailed in Table III. There are still some shortcomings that deserve attention. First, compared with the DAE model, the precision is slightly lower (95.53 vs. 96.92), indicating that there is space for improvement in the prediction accuracy of specific entities.
TABLE IV: Comparison Experimental results at Weibo.
@@
FLAT Li 2020 [2] - - 63.42
FLAT+BERT Li 2020 [2] - - 68.55
@@
Proposed method achieves an F1 score of 68.60, slightly higher than FLAT+BERT (68.55) and close to Glyce (67.60), demonstrating robustness in complex text processing on Weibo, shown as Table IV. Furthermore, proposed method performs well in recall (70.76), indicating its advantage in identifying more potential entities. However, compared with the precision of DAE (69.68), proposed method still has a certain gap in accuracy, which suggests that proposed model still has room for improvement in accurate prediction when dealing with noisy social media data.
IV-D2 Ablation Studies
Ablation Study in Table V evaluates the proposed structure through two experiments: (1) module removal to assess synergy, and (2) BERT replacement/removal to analyze word embedding impacts.
@@ IV-D2 Ablation Studies
The experimental results of comparing the base model (S.N.1) with the models that introduce different BERT variants (S.N.2 and S.N.3) show that RoBERTa outperforms BERT in terms of vocabulary vectorization.
@@
TABLE V: Ablation Studies on MSRA
@@ A H R BERT P R F1
8 (Ours) $\surd$ $\surd$ RoBERTa 94.58 93.98 94.28
IV-E Parameter Selection
@@
TABLE VI: Effects of different parameters on performance in RHEMA.
@@
As seen from the Table VI, too few attention heads with 16 or 32, may cause the model to be unable to fully capture feature representations, thereby limiting the performance of the model. When the number of attention heads increases to 64, the performance of the model reaches its peak. However, when the number of attention heads is further increased to 128), the performance of the model does not continue to improve, but instead shows a slight decline. This phenomenon may be due to the fact that too many attention heads increase the complexity of the model, thereby introducing the risk of overfitting.
IV-F Case Analysis